two parallel plates (each of area 200cm2) of a capacitor are given equal and opposite charge of magnitude 1.75microC.The space between the plates is filled with dielectric.If the electric field in the dielectric is 1.5*10^6 V/m what will be the magnitude of chare induced on each dielelectric plate
can anyone solve it fully upto ans

Grade:12

1 Answers

Susmita
5 years ago
Q=1.75microC=1.75×10-6C
A=200cm2=200×10-4 m2
o=8.854×10-12 F/m
Eo=Q/A€o=9.8×106V/m
Ep=1.5×106V/m
E=Eo-Ep=8.3×106 V/m
Eo=kE
or,k=Eo/E=1.18
Induced charge 
Q'=(k-1)Q/k=0.266 micro C